# Max-Rank: Efficient Multiple Testing for Conformal Prediction

### v3

@@ Abstract
Multiple hypothesis testing (MHT) commonly arises in various scientific fields, from genomics to psychology, where testing many hypotheses simultaneously increases the risk of Type-I errors. These errors can mislead scientific inquiry, rendering MHT corrections essential. In this paper, we address MHT within the context of conformal prediction, a flexible method for predictive uncertainty quantification. Some conformal prediction settings can require simultaneous testing, and positive dependencies among tests typically exist. We propose a novel correction named max-rank that leverages these dependencies, whilst ensuring that the joint Type-I error rate is efficiently controlled. Inspired by permutation-based corrections from Westfall and Young (1993), max-rank exploits rank order information to improve performance, and readily integrates with any conformal procedure. We demonstrate both its theoretical and empirical advantages over the common Bonferroni correction and its compatibility with conformal prediction, highlighting the potential to enhance predictive uncertainty quantification.
@@ 1 INTRODUCTION
Multiple hypothesis testing (MHT) occurs whenever we attempt to draw conclusions over multiple statistical hypothesis at once (Shaffer, 1995). Such scenarios frequently occur in various scientific domains, ranging from medicine and genomics (Ludbrook, 1994; Qu et al., 2010; Storey and Tibshirani, 2003) to economics and psychology (Blakesley et al., 2009; List et al., 2019). A crucial issue with MHT is that as the number of tested hypotheses increases, so does the likelihood of committing a Type-I error, i.e., erroneously rejecting a null in favour of significance. Such false positives may lead scientists astray by drawing incorrect conclusions and pursuing costly but statistically meaningless studies (Benjamini, 2010). Thus, a large body of research in statistics has preoccupied itself with multiple testing corrections to permit valid statistical inference even in such scenarios (Bender and Lange, 2001).
In this work, we investigate a particular instance of the multiple testing problem: predictive uncertainty quantification via conformal prediction (Shafer and Vovk, 2008; Vovk et al., 2005). Conformal prediction is an efficient way to produce statistically rigorous uncertainties for complex predictive models, such as deep neural networks (Angelopoulos et al., 2023). MHT issues can arise from the need to run multiple conformal procedures at once (e.g., in multi-target regression (Messoudi et al., 2021)), necessitating a correction. Moreover, these conformal procedures are known to exhibit positive dependency (Bates et al., 2022), rendering simple corrections such as Bonferroni inefficient a priori.
Exploiting the natural connections between conformal prediction and permutation testing, we propose a simple MHT correction that is able to exploit the positive dependency between hypotheses when apparent through rank ordering. Furthermore, it ensures strong Type-I error control on the family-wise error rate (FWER), i.e., the probability of a single false discovery. Our approach is closely connected to the permutation-based corrections of Westfall and Young (1993), and readily integrates with the conditions assumed by conformal prediction. To summarize, our contributions include:
  • Proposing an efficient correction called max-rank for MHT settings with positive dependency (Algo. 2) and establishing its connection to the correction procedures of Westfall and Young (1993) (§ 3.2).
  • Proving that max-rank is superior to the Bonferroni correction in terms of FWER control via multivariate copula arguments (§ 4, § 5).
  • Demonstrating both theoretically (§ 6.1) and empirically (§ 6.2) max-rank’s suitability for conformal prediction.
@@ 2 BACKGROUND & RELATED WORK
We begin by commenting on our multiple testing setting and forms of Type-I error control, with a closer look at the Bonferroni correction. We then thoroughly introduce the conformal prediction setting and its connection to hypothesis testing.
Multiple testing in parallel and in series. A useful distinction can be made between statistical testing in parallel and in series (Good, 1958). When testing in parallel, multiple hypotheses are examined simultaneously. Frequently, conclusions are drawn on the basis of the same underlying evidence such as a single common dataset (Vovk et al., 2022), making independence assumptions harder to justify. This is our setting of interest, and the one underlying conformal prediction and any arising MHT issues (see § 6). In contrast, testing in series suggests conclusions are drawn one after another, and the outcome of a test may determine whether and which subsequent hypotheses should be explored. For example, a clinical drug trial with multiple phases may only initiate a follow-up if the results are satisfactory during the first phase (Bauer, 1991).
Family-wise error rate control. We may distinguish two forms of controlling the Type-I error rate during testing: control in the strong sense via the family-wise error rate (FWER), and in the weaker sense via the false discovery rate (FDR, see § C.1). FWER control, the more stringent of the two, aims to limit the probability of committing at least one Type-I error across all considered hypotheses. If we test $m$ hypotheses and denote by $m_{\text{FP}}$ the number of falsely rejected null hypotheses (i.e., false positives), then FWER control at significance level $\alpha$ ensures that
@@
Such strict error control is appropriate in settings where false discoveries may have strong implications, such as in drug testing and clinical trials. Since conformal prediction is oftentimes motivated from a safety perspective by its provided coverage guarantees (see § 2.1) and can be applied to similar settings (e.g., Lei and Candès (2021); Candès et al. (2023); Yin et al. (2024)), FWER control is also an appropriate notion here.
@@
Multiple testing under dependency. Albeit MHT issues have been addressed as early as Tukey (1953), the research on valid and efficient corrections under dependency remains remarkably active. Westfall and Young (1993) suggested an early resampling-based procedure for capturing data dependencies, while recent years have seen further work on arbitrary dependencies (Causeur et al., 2009; Sun and Tony Cai, 2009; Leek and Storey, 2008; Liu and Xie, 2019; Finner et al., 2017; Gou and Tamhane, 2018; Chi et al., 2022) and p-value merging (Wilson, 2019; Vovk et al., 2022; Chen et al., 2023). Indeed, a remark by Chi et al. (2022) that “the gap between theory and practice is mildly unsettling” suggests these problems remain far from fully solved.
Multiple testing in conformal prediction. Conformal applications necessitating some form of multiple testing correction include multi-step time series (Stankeviciute et al., 2021; Sun and Yu, 2023; Xu et al., 2024) and trajectory estimation (Lindemann et al., 2023; Muthali et al., 2023); multi-target classification (Cauchois et al., 2021; Fisch et al., 2021) and regression (Messoudi et al., 2021, 2022; Feldman et al., 2023), including conformal object detection (Andéol et al. (2023); Mukama et al. (2024), see § 6.2); treatment analysis (Lei and Candès, 2021; Yin et al., 2024) and screening (Jin and Candès, 2023; Fisch et al., 2022), or even accelerated inference (Schuster et al., 2021). Throughout, employed corrections range from Bonferroni (Stankeviciute et al., 2021; Lindemann et al., 2023; Muthali et al., 2023; Andéol et al., 2023; Jin and Candès, 2023) to Copulas (Messoudi et al., 2021; Sun and Yu, 2023), Simes (Fisch et al., 2021) and other forms of sequential testing (Lei and Candès, 2021; Yin et al., 2024; Cauchois et al., 2021; Angelopoulos et al., 2021), showcasing that a variety of solutions have been considered.
2.1 Conformal prediction
@@
1:Input: Pre-trained predictor $\hat{f}$, calibration data $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim\mathcal{P}_{XY}$, test sample $(X_{n+1},Y_{n+1})\sim\mathcal{P}_{XY}$, miscoverage rate $\alpha\in(0,1)$.
@@
6:For the new test sample $(X_{n+1},Y_{n+1})$, a valid prediction set for $X_{n+1}$ is given by $\hat{C}(X_{n+1})=\{y\in\mathcal{Y}:\,s(\hat{f}(X_{n+1}),y)\leq\hat{Q}(1-\alpha;S)\}$. Validity refers to satisfying Eq. 2.
@@
Conformal prediction (CP) is a distribution-free, post-hoc, and finite-sample approach to predictive uncertainty quantification (Vovk et al., 2005; Shafer and Vovk, 2008). The obtained uncertainty around a predictor’s outputs is expressed in terms of prediction sets, and provides an associated guarantee on coverage (Eq. 2). The general algorithmic recipe for split CP (Papadopoulos et al., 2007), its most popular variant, is outlined in Algo. 1. Split CP requires a separate calibration dataset $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim\mathcal{P}_{XY}$ on which dissimilarity measures (so-called nonconformity scores) between predicted and true values are computed. For a new test sample $(X_{n+1},Y_{n+1})$, the prediction set is then constructed by including any predictions that align with the previously observed scores on $\mathcal{D}_{cal}$, as determined by a conformal quantile $\hat{Q}_{1-\alpha}$ that ensures a coverage rate of $(1-\alpha)$ for true values.
The approach is distribution-free since it does not impose any distributional assumptions. Instead, the key requirement is the exchangeability of $\mathcal{D}_{cal}\,\cup\,\{(X_{n+1},Y_{n+1})\}$, a relaxed i.i.d. assumption suggesting permutation invariance of the samples. It is post-hoc because it can be used with any ‘black-box’ pre-trained predictor without the need for training access. And it is finite-sample because the obtainable coverage guarantees do not rely on asymptotics. Specifically, if the conditions of CP are met, a probabilistic statement on containment of $Y_{n+1}$ in the prediction set $\hat{C}(X_{n+1})$ for some tolerated miscoverage rate $\alpha\in(0,1)$ is given as
@@
Importantly, the provided coverage guarantee is marginally valid, i.e., it only holds on average across test and calibration data and may thus fail to cover for any particular observation.
Conformal prediction as hypothesis testing. The CP procedure in Algo. 1 can also be interpreted from a hypothesis testing perspective (Shi et al., 2013; Bates et al., 2022; Vovk et al., 2005). Given a prediction set candidate value $y\in\mathcal{Y}$, a statistical hypothesis for the test sample is established as $H_{0}:Y_{n+1}=y,\,\,H_{1}:Y_{n+1}\neq y$, with test statistic $s_{n+1}=s(\hat{f}(X_{n+1}),y)$ equal to its nonconformity score. The set of calibration scores $S=\{s_{1},\dots,s_{n}\}$ provides an empirical null distribution to compute a p-value $\hat{P}_{n+1}(y;S)$ as
$\hat{P}_{n+1}(y;S)=|\{i=1,\dots,n:s_{i}\geq s_{n+1}\}|/(n+1),$ (3)
i.e., the fraction of samples which conform worse than $s_{n+1}$. A testing decision on not rejecting $H_{0}$ – that is, inclusion in the prediction set – is taken at significance level $\alpha$ as $\hat{P}_{n+1}(y;S)>\alpha$. This approach can be interpreted as permutation testing: we are assessing if the observed nonconformity for our candidate value does not stray too much from previous observations, and the sample $(X_{n+1},y)$ thus permutes with $D_{cal}$. We may then interpret running multiple conformal procedures at once as permutation testing in parallel, typically at identical significance level $\alpha$ across tests.
Equivalence of events. By the prediction set construction outlined in Algo. 1 and the above testing perspective, a useful equivalence of events for some set candidate value $y\in\mathcal{Y}$ can be established as
@@
Notably, the conformal quantile $\hat{Q}_{1-\alpha}$ [^1] We use $\hat{Q}_{1-\alpha}$ as a shortened notation for $\hat{Q}(1-\alpha;S)$. ensuring a fraction of $(1-\alpha)$ values below it is the equivalent of a permutation test at significance level $\alpha$. In both cases, the Type-I error is controlled at the same level $\alpha$. We further comment on the quantile’s properties in the following remarks:
@@
The smallest possible quantile $\hat{Q}_{1-\alpha}$ which maintains Type-I error control at level $\alpha$ is tight in terms of coverage, i.e., it provides the smallest and thus most efficient conformal prediction sets.
@@ 3 MULTIPLE TESTING CORRECTION VIA MAX-RANK
Notation. For consistency, we denote matrices as upper-case bolded ($\mathbf{M}$), vectors as lower-case bolded ($\mathbf{v}$), scalars without bolding ($s$). A multivariate random variable may also be upper-case bolded given the context ($\mathbf{X}$). An index set $\{1,\dots,n\}$ is abbreviated $[n]$ for any $n\geq 1$. Similarly, a sequence $x_{1},\dots,x_{n}$ is shortened to $x_{1:n}$. The rank of an element $x_{i}$ in a set $\{x_{1:n}\}$ of $n$ distinct elements is given by
@@
i.e., the size of the set of elements below $x_{i}$. Finally, the $l^{\infty}$-norm is given by $\lVert\mathbf{x}\rVert_{\infty}=\max_{1\leq i\leq n}|x_{i}|$ for a discrete vector $\mathbf{x}$ of size $n$.

### v4

@@ Abstract
Multiple hypothesis testing (MHT) frequently arises in scientific inquiries, and concurrent testing of multiple hypotheses inflates the risk of Type-I errors or false positives, rendering MHT corrections essential. This paper addresses MHT in the context of conformal prediction, a flexible framework for predictive uncertainty quantification. Some conformal applications give rise to simultaneous testing, and positive dependencies among tests typically exist. We introduce max-rank, a novel correction that exploits these dependencies whilst efficiently controlling the family-wise error rate. Inspired by existing permutation-based corrections, max-rank leverages rank order information to improve performance and integrates readily with any conformal procedure. We establish its theoretical and empirical advantages over the common Bonferroni correction and its compatibility with conformal prediction, highlighting the potential to strengthen predictive uncertainty estimates.
@@ 1 INTRODUCTION
Multiple hypothesis testing (MHT) occurs whenever we attempt to draw conclusions over multiple statistical hypotheses at once (Shaffer, 1995). Such scenarios regularly occur across different scientific domains ranging from medicine and genomics (Ludbrook, 1994; Qu et al., 2010; Storey and Tibshirani, 2003) to economics and psychology (Blakesley et al., 2009; List et al., 2019). A key issue is that as the number of tests grows, so does the risk of Type-I errors – falsely rejecting the null hypothesis – which can mislead scientific inquiry and waste resources on spurious or incorrect findings (Benjamini, 2010). To mitigate this, a vast body of statistical research has focused on multiple testing corrections that permit valid inference even in such cases (Bender and Lange, 2001).
In this work, we examine a specific instance of the multiple testing problem: predictive uncertainty quantification through conformal prediction (Shafer and Vovk, 2008; Vovk et al., 2005). Conformal prediction offers a statistically rigorous framework for uncertainty estimation in complex predictive models, such as deep neural networks (Angelopoulos et al., 2023). However, MHT challenges arise when multiple conformal procedures are run simultaneously (e.g., for multi-target regression (Messoudi et al., 2021)), necessitating a correction. Moreover, these procedures often exhibit positive dependencies (Bates et al., 2022), rendering standard approaches like Bonferroni inefficient a priori.
footnotetext: Corresponding author: $<$a.r.timans@uva.nl$>$footnotetext: ${}^{\text{\textdagger}}$ignificant contributions.
Leveraging the natural connections between conformal prediction and permutation testing, we introduce a simple MHT correction that exploits positive dependencies among hypotheses when evident through rank ordering. Our method ensures strong control of the family-wise error rate (FWER), i.e., the probability of at least one false discovery. Closely related to the permutation-based corrections of Westfall and Young (1993), it readily integrates with the conditions assumed by conformal prediction without imposing any additional assumptions. In summary, we contribute:
  • max-rank, an efficient correction for MHT settings with positive dependency (Algo. 2), and establish its connection to the multiple testing procedures of Westfall and Young (1993) (§ 3.2).
  • Prove that max-rank offers superior FWER control over the Bonferroni correction using multivariate copula arguments (§ 4, § 5).
  • Demonstrate both theoretically (§ 6.1) and empirically (§ 6.2) max-rank’s suitability for conformal prediction settings with multiple testing.
@@ 2 BACKGROUND & RELATED WORK
We first discuss our multiple testing setting and Type-I error control, with a closer look at the Bonferroni correction. Then, we introduce the conformal prediction setting and its connection to hypothesis testing.
Multiple testing in parallel and in series. A useful distinction can be made between statistical testing in parallel and in series (Good, 1958). When testing in parallel, multiple hypotheses are examined simultaneously. Frequently, conclusions are drawn on the basis of the same underlying evidence such as a single common dataset (Vovk et al., 2022), making independence assumptions harder to justify. This is our setting of interest, and the one underlying conformal prediction and any arising MHT issues (see § 6). In contrast, testing in series suggests conclusions are drawn one after another, and the outcome of a test may determine whether and which subsequent hypotheses should be explored. For example, a clinical drug trial with multiple phases may only initiate a follow-up if results are satisfactory during the first phase (Bauer, 1991).
Family-wise error rate control. We can further distinguish among different types of Type-I error control while testing: control in the strong sense via the FWER, and in weaker forms via, e.g., the false discovery rate (see § C.1). FWER control as the most stringent form aims to limit the probability of committing at least one Type-I error across all considered hypotheses. If we test $m$ hypotheses and denote by $m_{\text{FP}}$ the number of falsely rejected null hypotheses (i.e., false positives), then FWER control at significance level $\alpha$ ensures that
@@
This strict error control is appropriate in settings where false discoveries may have strong implications, such as in drug testing or clinical trials. Since conformal prediction is oftentimes motivated from a safety perspective by its provided coverage guarantees and is applicable to such settings (e.g., Lei and Candès (2021); Candès et al. (2023); Yin et al. (2024)), FWER control is also the appropriate notion here.
@@
Multiple testing under dependency. Albeit MHT issues have been addressed as early as Tukey (1953), the research on valid and efficient corrections under dependency remains remarkably active. Westfall and Young (1993) suggested an early resampling-based procedure for capturing data dependencies, while recent years have seen further work on arbitrary dependencies (Causeur et al., 2009; Sun and Tony Cai, 2009; Leek and Storey, 2008; Liu and Xie, 2019; Finner et al., 2017; Gou and Tamhane, 2018; Chi et al., 2025) and p-value merging (Wilson, 2019; Vovk et al., 2022; Chen et al., 2023). Indeed, a remark by Chi et al. (2025) that “the gap between theory and practice is mildly unsettling” suggests these questions remain far from fully solved.
Multiple testing in conformal prediction. Conformal applications necessitating some form of multiple testing correction are widespread, and include multi-step time series and trajectory estimation (Stankeviciute et al., 2021; Sun and Yu, 2023; Xu et al., 2024; Lindemann et al., 2023; Muthali et al., 2023); multi-target classification and regression (Cauchois et al., 2021; Fisch et al., 2021; Messoudi et al., 2021, 2022; Feldman et al., 2023), among them conformal object detection (Andéol et al. (2023); Mukama et al. (2024), see § 6.2); treatment analysis and screening (Lei and Candès, 2021; Yin et al., 2024; Jin and Candès, 2023; Fisch et al., 2022), and even as diverse as early-exiting and diffusion models (Schuster et al., 2021; Teneggi et al., 2023).
Throughout, employed MHT corrections frequently include Bonferroni (Stankeviciute et al., 2021; Lindemann et al., 2023; Muthali et al., 2023; Andéol et al., 2023; Jin and Candès, 2023), but can range from Copulas (Messoudi et al., 2021; Sun and Yu, 2023) to Simes (Fisch et al., 2021) and other forms of sequential testing (Lei and Candès, 2021; Yin et al., 2024; Cauchois et al., 2021; Angelopoulos et al., 2021), showcasing a variety of considered solutions. Naturally, the application-specific test design (in parallel vs. in series) largely governs the suitability of employing particular corrections.
@@
1:Input: Pre-trained predictor $\hat{f}$, calibration samples $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim\mathcal{P}_{XY}$, test sample $(X_{n+1},Y_{n+1})\sim\mathcal{P}_{XY}$, miscoverage rate $\alpha\in(0,1)$.
@@
6:For a new test sample $(X_{n+1},Y_{n+1})$, a valid prediction set satisfying Eq. 2 is given by $\hat{C}(X_{n+1})=\{y\in\mathcal{Y}:\,s(\hat{f}(X_{n+1}),\,y)\leq\hat{Q}(1-\alpha;S)\}$.
@@
Conformal prediction. Conformal prediction (CP) offers a distribution-free, post-hoc, and finite-sample approach to predictive uncertainty quantification (Vovk et al., 2005; Shafer and Vovk, 2008). The obtained uncertainty around a predictor’s output is expressed as a prediction set, and provides an associated guarantee on coverage (Eq. 2). The general algorithmic recipe for split CP (Papadopoulos et al., 2007), its most popular variant, is outlined in Algo. 1. Split CP requires a separate calibration dataset $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim\mathcal{P}_{XY}$ for which dissimilarity values (so-called nonconformity scores) between predicted and true outputs are computed. Given a new test sample $(X_{n+1},Y_{n+1})\sim\mathcal{P}_{XY}$, a prediction set is then constructed by including any predictions that align with the previously observed scores on $\mathcal{D}_{cal}$, as governed by a conformal quantile $\hat{Q}_{1-\alpha}$ ensuring a coverage rate of $(1-\alpha)$ for true values. The approach is distribution-free since it does not impose any distributional assumptions. Instead, a key requirement is the exchangeability of $\mathcal{D}_{cal}\,\cup\,\{(X_{n+1},Y_{n+1})\}$, a relaxed i.i.d. assumption suggesting permutation invariance of samples. It is post-hoc because it can be used with any ‘black-box’ pre-trained predictor without the need for training or internal access. And it is finite-sample because the obtainable coverage guarantee does not rely on asymptotics. Specifically, if CP’s conditions are met, a probabilistic statement on inclusion of $Y_{n+1}$ in the prediction set $\hat{C}(X_{n+1})$ for a tolerated miscoverage rate $\alpha\in(0,1)$ is given as
@@
Importantly, the provided coverage guarantee is (i) marginally valid, i.e., it only holds on average across test and calibration data and may thus fail to cover for any particular observation; and (ii) fails to appropriately cover if exchangeability is violated, such as under strong distribution shifts (Barber et al., 2023).
Conformal prediction as hypothesis testing. The CP procedure in Algo. 1 can also be interpreted from a hypothesis testing perspective (Shi et al., 2013; Bates et al., 2022; Vovk et al., 2005). Given a candidate value $y\in\mathcal{Y}$, a statistical test for its inclusion in the prediction set can be formalized by the hypothesis pair $H_{0}:Y_{n+1}=y,\,\,H_{1}:Y_{n+1}\neq y$, with test statistic $s_{n+1}=s(\hat{f}(X_{n+1}),y)$ equating its nonconformity score. The set of calibration scores $S=\{s_{1},\dots,s_{n}\}$ provides an empirical null distribution to compute a rank-based p-value $\hat{P}_{n+1}(y;S)$ as
$\hat{P}_{n+1}(y;S)=\frac{|\{i=1,\dots,n+1:s_{i}\geq s_{n+1}\}|}{n+1},$ (3)
i.e., the fraction of samples which conform worse than $s_{n+1}$. A testing decision on not rejecting $H_{0}$ – that is, inclusion in the prediction set – is taken at significance level $\alpha$ as $\hat{P}_{n+1}(y;S)>\alpha$. The formulated testing approach translates to an instance of permutation testing: we are assessing if the observed nonconformity for candidate value $y$ does not stray overly far from previous observations, and the sample $(X_{n+1},y)$ thus permutes with $D_{cal}$. Subsequently, we interpret running multiple conformal procedures at once as permutation testing in parallel, and at identical level $\alpha$ across tests.
Equivalence of events. By the prediction set construction outlined in Algo. 1 and the above testing perspective, a useful equivalence of events for a prediction set candidate value $y$ can be established as
@@
Notably, determining set inclusion via the conformal quantile $\hat{Q}(1-\alpha;S)$ (abbreviated to $\hat{Q}_{1-\alpha}$) which ensures a coverage rate of $(1-\alpha)$ matches a permutation test using $\hat{P}_{n+1}(y;S)$ at significance level $\alpha$. In both cases, Type-I error control at the same rate $\alpha$ is achieved. We further comment on the quantile’s properties in the following remarks:
@@
The smallest possible quantile $\hat{Q}_{1-\alpha}$ which maintains Type-I error control at level $\alpha$ is tight in terms of coverage, i.e., it provides the smallest and thus most efficient conformal prediction sets. We may loosely interpret this as the matching permutation test ensuring highest statistical power.
@@ 3 MULTIPLE TESTING CORRECTION VIA MAX-RANK
Notation. Matrices are denoted as upper-case bold ($\mathbf{M}$), vectors as lower-case bold ($\mathbf{v}$), scalars without bolding ($s$). A multivariate random variable may also be upper-case bold given the context ($\mathbf{X}$). An index set $\{1,\dots,n\}$ is abbreviated $[n]$ for any $n\geq 1$. Similarly, a sequence $x_{1},\dots,x_{n}$ is shortened to $x_{1:n}$. The rank of element $x_{i}$ in a set $\{x_{1:n}\}$ of $n$ distinct elements (potentially by introducing jitter) is given by
@@
i.e., the size of the set of elements below $x_{i}$. Finally, the $l^{\infty}$-norm is given by $\lVert\mathbf{x}\rVert_{\infty}=\max_{1\leq i\leq n}|x_{i}|$ for a discrete vector $\mathbf{x}$ of size $n$. See also Table 2.